# Solvability of the Korteweg-de Vries equation under meromorphic initial conditions by quadrature

### v1

@@ A.1. Fundamental results
$v_{xx}=r(x)v,$ (A.1)
where $r(x)\in\mathbb{C}(x)\setminus\mathbb{C}$. Let $G$ be the differential Galois group of (A.1) over $\mathbb{C}(x)$. We easily see that $G$ is an algebraic subgroup of $\mathrm{SL}(2,\mathbb{C})$ (see, e.g., Chapter 6 of [12]). We have the following four cases for (A.1) (see [10, 13] for a proof).
Proposition A.1 (Kaplansky [12], Kovacic[13]).
@@
If one of cases (a)-(c) of Proposition A.1 occurs, then the connected identity component $G^{0}$ of $G$ is solvable and Eq. (A.1) is solvable. On the other hand, if cases (a)-(c) do not occur, then $G=G^{0}=\mathrm{SL}(2,\mathbb{C})$, i.e., Eq. (A.1) is not integrable in the sence of differential Galois theory.
Let $r(x)=r_{1}(x)/r_{2}(x)$, where $r_{1}(x),r_{2}(x)\in\mathbb{C}[z]$ are relatively prime. Poles of $r(x)$ are zeros of $r_{2}(x)$ and their orders are the multiplicities of the zeros of $r_{2}(x)$. Let $\mathscr{S}\subset\mathbb{C}$ denote the set of all poles of $r(x)$. By the order of $r(x)$ at $\infty$, we also mean the order of $\infty$ as a zero of $r(x)$, i.e., $\deg r_{2}(x)-\deg r_{1}(x)$. We have the following necessary conditions for cases (a)-(c) in Proposition A.1 to occur (see [10, 13] for a proof).
Proposition A.2 (Kovacic [13]).
The following conditions are necessary for the corresponding cases in Proposition A.1 to occur::\,::
@@
Note that cases (a), (b) and (c) of Proposition A.1 may not occur even though conditions (a), (b) and (c) hold, respectively. In particular, Proposition A.2 only states fewer conditions for case (c) than the original version in [13] since they are enough for our application in Section 5. Kovacic [13] also gave an algorithm for finding a “closed-form” solution and detecting whether cases (a)-(c) in Proposition A.1 really occur or not under conditions (a), (b) and (c) of Proposition A.1, respectively. In the following we describe necessary parts of his algorithm for cases (a) and (b) in the setting required in Section 5: The order of $r(x)$ at $\infty$ is zero. See [10, 13] for the full version of his algorithm and its proof.
A.2. Case (a)
We assume that condition (a) in Proposition A.2 holds as well as the order of $r(x)$ at $\infty$ is zero. When $g(x)$ has a pole of order $\ell\geq 2$ at $x=s$ and is expanded in a Laurent series
@@ A.2. Case (a)
$d_{c}=\kappa_{\infty}^{c(\infty)}-\sum_{s\in\mathscr{S}}\kappa_{s}^{c(s)}.$ (A.2)
If $d_{c}$ is not a non-negative integer for any sequence $c$, then case (a) of Proposition A.1 does not occur. We also have a further recipe to determine whether it occurs or not when $d_{c}$ is a non-negative integer, although it is not given here since it is not necessary for our purpose. The reader who is interested in it should consult [10, 13].
A.3. Case (b)
We assume that condition (b) in Proposition A.2 holds as well as the order of $r(x)$ at $\infty$ is zero. For $s\in\mathscr{S}\cup\{\infty\}$ we define a set $E_{s}\subset\mathbb{Z}$ as follows:
@@ A.3. Case (b)
$d_{e}=-\tfrac{1}{2}\sum_{s\in\mathscr{S}}e_{s}.$ (A.3)
If $d_{e}$ is not a non-negative integer for any $e$, then case (b) of Proposition A.1 does not occur.
@@
$\theta(x)=\tfrac{1}{2}\sum_{s\in\mathscr{S}}\frac{e_{s}}{x-s}$ (A.4)
for $e\in E$. By (A.3) and (A.4) we see that the Laurent series of $\theta(x)$ at $\infty$ is given by
@@
$+(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}-4r(x)\theta(x)-2r^{\prime}(x))P=0,$ (A.5)
then case (b) of Proposition A.1 occurs and
$v(x)=\exp\left(\int\tfrac{1}{2}\left(-\hat{\theta}(x)\pm\sqrt{4r(x)-\hat{\theta}(x)^{2}-2\hat{\theta}^{\prime}(x)}\right)\mathrm{d}x\right)$ (A.6)
are solutions to (A.1), where $\hat{\theta}(x)=\theta(x)+P^{\prime}(x)/P(x)$. If such a polynomial is not found for any sequence $e$, then it does not occur.

### v2

@@ B.1. Fundamental results
$v_{xx}=r(x)v,$ (B.1)
where $r(x)\in\mathbb{C}(x)\setminus\mathbb{C}$. Let $G$ be the differential Galois group of (B.1) over $\mathbb{C}(x)$. We easily see that $G$ is an algebraic subgroup of $\mathrm{SL}(2,\mathbb{C})$ (see, e.g., Chapter 6 of [12]). We have the following four cases for (B.1) (see [10, 13] for a proof).
Proposition B.1 (Kaplansky [12], Kovacic[13]).
@@
If one of cases (a)-(c) of Proposition B.1 occurs, then the connected identity component $G^{0}$ of $G$ is solvable and Eq. (B.1) is solvable. On the other hand, if cases (a)-(c) do not occur, then $G=G^{0}=\mathrm{SL}(2,\mathbb{C})$, i.e., Eq. (B.1) is not integrable in the sence of differential Galois theory.
Let $r(x)=r_{1}(x)/r_{2}(x)$, where $r_{1}(x),r_{2}(x)\in\mathbb{C}[z]$ are relatively prime. Poles of $r(x)$ are zeros of $r_{2}(x)$ and their orders are the multiplicities of the zeros of $r_{2}(x)$. Let $\mathscr{S}\subset\mathbb{C}$ denote the set of all poles of $r(x)$. By the order of $r(x)$ at $\infty$, we also mean the order of $\infty$ as a zero of $r(x)$, i.e., $\deg r_{2}(x)-\deg r_{1}(x)$. We have the following necessary conditions for cases (a)-(c) in Proposition B.1 to occur (see [10, 13] for a proof).
Proposition B.2 (Kovacic [13]).
The following conditions are necessary for the corresponding cases in Proposition B.1 to occur::\,::
@@
Note that cases (a), (b) and (c) of Proposition B.1 may not occur even though conditions (a), (b) and (c) hold, respectively. In particular, Proposition B.2 only states fewer conditions for case (c) than the original version in [13] since they are enough for our application in Section 5. Kovacic [13] also gave an algorithm for finding a “closed-form” solution and detecting whether cases (a)-(c) in Proposition B.1 really occur or not under conditions (a), (b) and (c) of Proposition B.1, respectively. In the following we describe necessary parts of his algorithm for cases (a) and (b) in the setting required in Section 5: The order of $r(x)$ at $\infty$ is zero. See [10, 13] for the full version of his algorithm and its proof.
B.2. Case (a)
We assume that condition (a) in Proposition B.2 holds as well as the order of $r(x)$ at $\infty$ is zero. When $g(x)$ has a pole of order $\ell\geq 2$ at $x=s$ and is expanded in a Laurent series
@@ B.2. Case (a)
$d_{c}=\kappa_{\infty}^{c(\infty)}-\sum_{s\in\mathscr{S}}\kappa_{s}^{c(s)}.$ (B.2)
If $d_{c}$ is not a non-negative integer for any sequence $c$, then case (a) of Proposition B.1 does not occur. We also have a further recipe to determine whether it occurs or not when $d_{c}$ is a non-negative integer, although it is not given here since it is not necessary for our purpose. The reader who is interested in it should consult [10, 13].
B.3. Case (b)
We assume that condition (b) in Proposition B.2 holds as well as the order of $r(x)$ at $\infty$ is zero. For $s\in\mathscr{S}\cup\{\infty\}$ we define a set $E_{s}\subset\mathbb{Z}$ as follows:
@@ B.3. Case (b)
$d_{e}=-\tfrac{1}{2}\sum_{s\in\mathscr{S}}e_{s}.$ (B.3)
If $d_{e}$ is not a non-negative integer for any $e$, then case (b) of Proposition B.1 does not occur.
@@
$\theta(x)=\tfrac{1}{2}\sum_{s\in\mathscr{S}}\frac{e_{s}}{x-s}$ (B.4)
for $e\in E$. By (B.3) and (B.4) we see that the Laurent series of $\theta(x)$ at $\infty$ is given by
@@
$+(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}-4r(x)\theta(x)-2r^{\prime}(x))P=0,$ (B.5)
then case (b) of Proposition B.1 occurs and
$v(x)=\exp\left(\int\tfrac{1}{2}\left(-\hat{\theta}(x)\pm\sqrt{4r(x)-\hat{\theta}(x)^{2}-2\hat{\theta}^{\prime}(x)}\right)\mathrm{d}x\right)$ (B.6)
are solutions to (B.1), where $\hat{\theta}(x)=\theta(x)+P^{\prime}(x)/P(x)$. If such a polynomial is not found for any sequence $e$, then it does not occur.